# Text2Earth: Unlocking Text-driven Remote Sensing Image Generation with a Global-Scale Dataset and a Foundation Model

### v1

@@ Abstract
Recently, generative foundation models have significantly advanced large-scale text-driven natural image generation and have become a prominent research trend across various vertical domains. However, in the remote sensing field, there is still a lack of research on large-scale text-to-image (text2image) generation technology. Existing remote sensing image-text datasets are small in scale and confined to specific geographic areas and scene types. Besides, existing text2image methods have struggled to achieve global-scale, multi-resolution controllable, and unbounded image generation. To address these challenges, this paper presents two key contributions: the Git-10M dataset and the Text2Earth foundation model. Git-10M is a global-scale image-text dataset comprising 10 million image-text pairs, 5 times larger than the previous largest one. The dataset contains essential resolution information and covers a wide range of geographic scenes and contains essential geospatial metadata, significantly surpassing existing datasets in both size and diversity. Building on Git-10M, we propose Text2Earth, a 1.3 billion parameter generative foundation model based on the diffusion framework to model global-scale remote sensing scenes. Text2Earth integrates a resolution guidance mechanism, enabling users to specify image resolutions. A dynamic condition adaptation strategy is proposed for training and inference to improve image generation quality. Text2Earth not only excels in zero-shot text2image generation but also demonstrates robust generalization and flexibility across multiple tasks, including unbounded scene construction, image editing, and cross-modal image generation. This robust capability surpasses previous models restricted to the basic fixed size and limited scene types. On the previous text2image benchmark dataset, Text2Earth outperfoms previous models with a significant improvement of +26.23 FID and +20.95% Zero-shot Cls-OA metric. Our project page is https://chen-yang-liu.github.io/Text2Earth/
@@ I Introduction
Recently, generative foundation models have significantly advanced large-scale text-driven natural image generation and have become a prominent research trend across various vertical domains [1, 2, 3], including medical imaging, autonomous driving, and virtual reality. These foundation models have demonstrated impressive image generation capabilities from large-scale image-text datasets, enabling them to produce large amounts of high-quality images. However, in the remote sensing field, there is still a lack of research on the large-scale text-to-image (text2image) generation technology based on foundation models [4, 5, 6]. This research holds considerable significance and application value, particularly in areas such as imaging simulation, virtual remote sensing scene construction, and data augmentation [7, 8, 9].
@@
In this paper, we aim to advance remote sensing text2image generation towards global-scale scene generation, multi-resolution controllability, and unbounded large-size image synthesis through two primary contributions: a large-scale dataset and a powerful generative foundation model. To overcome the limitations of existing datasets, we developed the Git-10M dataset, a Global-scale image-text dataset comprising 10 million image-text pairs, which is 5 times larger than the previous largest dataset. Surpassing previous datasets, Git-10M encompasses a diverse range of global geographical scenes, including cities, forests, and mountains, while also containing rich metadata such as image resolution and geographic location. This comprehensive diversity empowers the model trained on Git-10M to generate realistic and global-scale images across various geographic scenes.
@@ II-A Generative Foundation Models in the Computer Vision
Generative foundation models (GFMs) have become increasingly influential in the field of computer vision, demonstrating remarkable advancements in the generation and transformation of visual data [3, 23, 24, 25]. These models, which are based on large-scale pre-training, are designed to capture a broad range of visual concepts and structures from vast datasets, making them versatile for numerous downstream tasks. Current generative models mainly focus on text2image generation. These models are typically built on three architectures: Generative Adversarial Networks (GANs), Autoregressive Transformers, and Diffusion models.
@@ III Global-scale image-text Dataset
The Git-10M dataset is a global-scale remote sensing image-text pair dataset, consisting of 10 million image-text pairs with geographical locations and resolution information. This section will detail the construction process of the dataset and conduct a systematic analysis.
@@ III-A Image Collection and Preprocessing
As shown in Fig. 2, the images in the Git-10M dataset are sourced from multiple publicly available datasets and manually collected global remote sensing imagery from Google Earth. Public datasets, including Million-AID [74], GeoPile [75], SSL4EO-S12 [76], SkyScript [77], DIOR [78], and RSICB [79], provide high-quality remote sensing images. These datasets primarily focus on scene classification tasks. During the collection process, we retained the scene category labels for each image to enable more precise semantic descriptions during the subsequent text annotation phase. These diverse data sources significantly enhance the richness of the Git-10M dataset.
@@
After completing the image collection, we conducted stringent filtering and processing. First, duplicate or redundant ocean scenes were removed through manual screening to maintain diversity in geographic distribution. Additionally, a subset of images exhibited issues with visual quality, such as noise and artifact, which could negatively impact the training of image generation models. To address this, an image enhancement model was trained on a private high-quality remote sensing dataset and applied to all collected images, significantly improving the overall image quality of the dataset.
@@
Figure 3: The diverse geospatial distribution of the Git-10M dataset, spans multiple continents and geographical regions, covering various typical scenes such as urban areas, forests, mountains, and deserts.
@@ III-C Dataset Analysis
Figure 5: The quality score of images before and after enhancement processing. The results demonstrate a significant improvement after enhancement.
@@
    Geographical Coverage: We performed a statistical analysis of the geographical distribution of images in the Git-10M dataset. As shown in Fig. 3, Git-10M spans multiple continents and geographical regions, covering various typical scenes such as urban areas, forests, mountains, deserts, and more. The wide geographical coverage ensures that the dataset can support the generation of real-world remote sensing images across different regions, natural features, and diverse scenes.
@@
Figure 6: Left: the word cloud of the texts in the Git-10M dataset. Right: the distribution of text lengths in the Git-10M dataset shows that each textual description averages approximately 52 words, with the entire dataset comprising over 10.5 million text samples and more than 5.5 billion words.
@@ IV-A Structure of Text2Earth Model
Our Text2Earth can generate entirely new remote sensing images consistent with the provided text and resolution or perform local editing on existing images while preserving the original structure. Users can input a white mask to specify the image region for generating visual content, which can either encompass the entire image or focus on a specific localized area.
@@ IV-A1 Image Compression Encoding
Figure 7: The structure of the Text2Earth model equipped with 1.3 billion parameters. Text2Earth can generate entirely new images consistent with the provided text or perform local editing on existing images while preserving the original structure. Users can input a white mask to specify the image region for generating visual content, which can either encompass the entire image or focus on a specific localized area.
@@ V-A2 RSICD Dataset
RSICD dataset is a widely used benchmark dataset for remote sensing text2image generation. It contains 10,921 remote sensing images and corresponding text annotations. The dataset contains 30 types of common ground scenes, and the spatial resolution of images is not unique. This dataset was employed to evaluate our model’s adaptation to the small specific scene dataset.
To further explore the model’s capacity for multimodal image generation, we synthesized a multimodal dataset based on the RSICD dataset. RGB images in RSICD were transformed into various modalities as follows:
@@ V-C Evaluation Metrics
Following previous studies [7, 71], we also employ the Zero-Shot classification Overall Accuracy (Cls-OA) metric to evaluate the semantic alignment between generated images and their textual descriptions. A classification model (e.g., ResNet-18) is trained on generated images using text descriptions from the test set. This model is then used for zero-shot classification on the real test set without prior exposure to them during training. The OA metric thus measures the semantic coherence and relevance of the generated images to the textual prompts.
@@
Method FID Score $\downarrow$ Zero-Shot Cls-OA $\uparrow$
Attn-GAN [35] 95.81 32.56%
DAE-GAN [83] 93.15 29.74%
DF-GAN [27] 109.41 51.99%
Lafite [29] 74.11 49.37%
DALL-E [41] 191.93 28.59%
@@
CRS-Diff [71] 50.72 69.31%
Text2Earth (Ours) 24.49 90.26%
@@
Figure 9: Text2Earth can produce remote sensing images with different resolutions based on user specifications, from detailed close-ups to wide-area coverage.
@@ V-D Zero-Shot Text2Image Generation
Text2Earth also demonstrates remarkable robustness in generating realistic images, even in cases where key input conditions—such as text or resolution—are missing. For instance, the model can generate forest images at various scales when provided with text like “There is a dense forest” without a specified resolution. Similarly, when only the resolution is provided without specific textual descriptions, the model can still generate resolution-specific images with diverse scenes such as a forest or an urban area. This ability highlights the model’s robustness in handling incomplete or missing input data, making it adaptable for real-world applications where input conditions may be partial. This ability benefits from our proposed dynamic condition adaptation strategy described in Section IV-B.
@@ V-E Remote Sensing Image Editing
Importantly, Text2Earth ensures that these modifications are seamlessly integrated with the surrounding areas, maintaining continuity and coherence. This makes it an ideal tool for customized remote sensing image editing, catering to diverse application needs such as urban planning and landscape management.
@@ V-G1 Text-Driven Multi-Modal Image Generation
We fine-tuned Text2Earth using LoRA on the RSICD dataset to facilitate text-driven multi-modal image generation tasks, such as Text2SAR, Text2NIR, and Text2PAN. The results are illustrated in Fig. 13, with quantitative evaluations presented in Table IV. The experiments demonstrate that Text2Earth can generate multi-modal images with high quality and semantic consistency. For instance, in the generated NIR images, green vegetation areas exhibit high pixel values, aligning with the physical imaging principles of NIR, where green vegetation reflects strongly in the near-infrared spectrum. These results underscore Text2Earth’s ability to effectively transfer its general knowledge to multi-modal remote sensing image generation tasks.
@@ V-G2 Image-to-Image Translation
VI Discussion and Future Work

### v2

@@ Abstract
Recently, generative foundation models have significantly advanced large-scale text-driven natural image generation and have become a prominent research trend across various vertical domains. However, in the remote sensing field, there is still a lack of research on large-scale text-to-image (text2image) generation technology. Existing remote sensing image-text datasets are small in scale and confined to specific geographic areas and scene types. Besides, existing text2image methods have struggled to achieve global-scale, multi-resolution controllable, and unbounded image generation. To address these challenges, this paper presents two key contributions: the Git-10M dataset and the Text2Earth foundation model. Git-10M is a global-scale image-text dataset comprising 10.5 million image-text pairs, 5 times larger than the previous largest one. The dataset contains essential resolution information and covers a wide range of geographic scenes and contains essential geospatial metadata, significantly surpassing existing datasets in both size and diversity. Building on Git-10M, we propose Text2Earth, a 1.3 billion parameter generative foundation model based on the diffusion framework to model global-scale remote sensing scenes. Text2Earth integrates a resolution guidance mechanism, enabling users to specify image resolutions. A dynamic condition adaptation strategy is proposed for training and inference to improve image generation quality. Text2Earth not only excels in zero-shot text2image generation but also demonstrates robust generalization and flexibility across multiple tasks, including unbounded scene construction, image editing, and cross-modal image generation. This robust capability surpasses previous models restricted to the basic fixed size and limited scene types. On the previous text2image benchmark dataset, Text2Earth outperfoms previous models with a significant improvement of +26.23 FID and +20.95% Zero-shot Cls-OA metric. Our project page is https://chen-yang-liu.github.io/Text2Earth/
@@ I Introduction
Recently, generative foundation models have significantly advanced large-scale text-driven natural image generation and have become a prominent research trend across various vertical domains [1, 2, 3], including medical imaging, autonomous driving, and virtual reality. These foundation models have demonstrated impressive image generation capabilities from large-scale image-text datasets, enabling them to produce large amounts of high-quality images. However, in the remote sensing field, there is still a lack of research on the large-scale text-to-image (text2image) generation technology based on foundation models [4, 5, 6, 7]. This research holds considerable significance and application value, particularly in areas such as imaging simulation, virtual remote sensing scene construction, and data augmentation [8, 9, 10].
@@
In this paper, we aim to advance remote sensing text2image generation towards global-scale scene generation, multi-resolution controllability, and unbounded large-size image synthesis through two primary contributions: a large-scale dataset and a powerful generative foundation model. To overcome the limitations of existing datasets, we developed the Git-10M dataset, a Global-scale image-text dataset comprising 10.5 million image-text pairs, which is 5 times larger than the previous largest dataset. Surpassing previous datasets, Git-10M encompasses a diverse range of global geographical scenes, including cities, forests, and mountains, while also containing rich metadata such as image resolution and geographic location. This comprehensive diversity empowers the model trained on Git-10M to generate realistic and global-scale images across various geographic scenes.
@@ II-A Generative Foundation Models in the Computer Vision
Generative foundation models (GFMs) have become increasingly influential in the field of computer vision, demonstrating remarkable advancements in the generation and transformation of visual data [3, 24, 25, 26, 27]. These models, which are based on large-scale pre-training, are designed to capture a broad range of visual concepts and structures from vast datasets, making them versatile for numerous downstream tasks. Current generative models mainly focus on text2image generation. These models are typically built on three architectures: Generative Adversarial Networks (GANs), Autoregressive Transformers, and Diffusion models.
@@ III Global-scale image-text Dataset
The Git-10M dataset is a global-scale remote sensing image-text pair dataset, consisting of 10.5 million image-text pairs with geographical locations and resolution information. This section will detail the construction process of the dataset and conduct a systematic analysis.
@@ III-A Image Collection and Preprocessing
As shown in Fig. 2, the images in the Git-10M dataset are sourced from multiple publicly available datasets and manually collected global remote sensing imagery from Google Earth. The public datasets, including Million-AID [76], GeoPile [77], SSL4EO-S12 [78], SkyScript [79], DIOR [80], and RSICB [81], provide high-quality remote sensing images. These datasets primarily focus on scene classification tasks. During the collection process, we retained the scene category labels for each image to enable more precise semantic descriptions during the subsequent text annotation phase. These diverse data sources significantly enhance the richness of the Git-10M dataset.
@@
After completing the image collection, we conducted stringent filtering and processing. First, duplicate or redundant ocean scenes were removed through manual screening to maintain diversity in geographic distribution. Additionally, a subset of images exhibited issues with visual quality, such as noise and artifact, which could negatively impact the training of image generation models. To address this, an image enhancement model was trained on a private high-quality remote sensing dataset and applied to all collected images, significantly improving the overall image quality of the dataset. During the training of the model, we simulate various image degradation processes, such as blurring, noise addition, and compression, to create paired low-quality and high-quality images. We train the model using these paired images to learn the mapping from degraded images to their high-quality counterparts. This enhancement process helps to standardize image quality across the Git-10M dataset, making it more suitable for high-quality generative modeling. We will also release the enhancement model at https://github.com/Chen-Yang-Liu/Text2Earth
@@
Figure 3: The diverse geospatial distribution of the Git-10M dataset. The yellow pixels represent the geographic locations where remote sensing images in Git-10M were sampled. The distribution shows that our dataset covers multiple continents and geographical regions, covering various typical scenes such as urban areas, forests, mountains, and deserts.
@@ III-C Dataset Analysis
    Geographical Coverage: We performed a statistical analysis of the geographical distribution of images in the Git-10M dataset. As shown in Fig. 3, Git-10M spans multiple continents and geographical regions, covering various typical scenes such as urban areas, forests, mountains, deserts, and more. The wide geographical coverage ensures that the dataset can support the generation of real-world remote sensing images across different regions, natural features, and diverse scenes. Besides, as stated in the Section III-A, some images of our Git-10M dataset are collected from several public scene classification datasets, which provide explicit scene labels. The integration of these datasets ensures that Git-10M covers a wide variety of well-defined remote sensing scene types. For example, the AID dataset contains 30 typical scene categories, while the RSICB dataset contains 45 categories.
@@
    Figure 5: The quality score of images before and after enhancement processing for our Git-10M dataset. The results demonstrate a significant improvement after enhancement. An example is shown on the right.
@@
Figure 6: Text Analysis. Top: the word cloud of the texts in the Git-10M dataset. Bottom: the distribution of text lengths in the Git-10M dataset shows that each textual description averages approximately 52 words, with the entire dataset comprising over 10.5 million text samples and more than 5.5 billion words.
@@ IV-A Structure of Text2Earth Model
Our Text2Earth can generate entirely new remote sensing images consistent with the provided text and resolution or perform local editing on existing images while preserving the original structure. Users can input a white mask to specify the image region for generating visual content, which can either encompass the entire image or focus on a specific area.
@@ IV-A1 Image Compression Encoding
Figure 7: The structure of the Text2Earth model equipped with 1.3 billion parameters. Text2Earth can generate entirely new images consistent with the provided text or perform local editing on existing images while preserving the original structure. Users can input a white mask to specify the image region for generating visual content, which can either encompass the entire image or focus on a specific area.
@@ V-A2 RSICD Dataset
RSICD dataset is a widely used benchmark dataset for remote sensing text2image generation. It contains 10,921 remote sensing images and corresponding text annotations. The dataset contains 30 types of common ground scenes, and the spatial resolution of images is not unique. This dataset was employed to evaluate our model’s adaptation to the small specific scene dataset. To further explore the multimodal image generation, we extend the RSICD dataset to a multimodal dataset. RGB images in the RSICD dataset were transformed into various modalities as follows:
@@ V-C Evaluation Metrics
Following previous studies [8, 73], we also employ the Zero-Shot classification Overall Accuracy (Cls-OA) metric to evaluate the semantic alignment between generated images and their textual descriptions. Specifically, a classification model (i.e., ResNet-18) is trained on generated images using text descriptions from the test set. This model is then used for zero-shot classification on the real test set without prior exposure to them during training. The OA metric thus measures the semantic coherence and relevance of the generated images to the textual prompts.
Furthermore, to measure the semantic similarity between text and generated images, we also used the CLIP score, which is calculated as follows:
CLIP $\text{CLIP Score}=\frac{1}{N}\sum_{i=1}^{N}\cos\left(E_{\text{image}}(I_{i}),E_{\text{text}}(T_{i})\right)\times 100$
where $I_{i}$ denotes the $i$-th generated image, $T_{i}$ denotes the corresponding input text, $E_{\text{image}}$ and $E_{\text{text}}$ represent the image encoder and text encoder of a pretrained CLIP model, respectively, and $\cos(\cdot)$ denotes the cosine similarity between the two embedding vectors. The final CLIP score is the average cosine similarity across all $N$ text-image pairs, reflecting the overall semantic alignment quality.
@@
Method FID $\downarrow$ Zero-Shot Cls-OA $\uparrow$ CLIP Score $\uparrow$
Attn-GAN [37] 95.81 32.56% 20.19
DAE-GAN [85] 93.15 29.74% 19.69
DF-GAN [29] 109.41 51.99% 19.76
Lafite [31] 74.11 49.37% 22.52
DALL-E [43] 191.93 28.59% 20.13
@@
CRS-Diff [73] 50.72 69.31% 20.33
Text2Earth (Ours) 24.49 90.26% 25.62
@@
Figure 9: Generated images with different resolutions solely by specifying the resolution condition, without any descriptive text.. Text2Earth can generate images reflecting a range of spatial resolutions—from high-resolution close-up views that capture fine details to lower-resolution images that cover larger areas. For example, in the generated images of mountainous regions, higher-resolution images exhibit detailed terrain features, while lower-resolution images depict broader landscape coverage, which aligns with real-world spatial resolution characteristics.
Figure 10: Resolution-conditioned image generation with same text prompts.
@@ V-D Zero-Shot Text2Image Generation
Text2Earth also demonstrates remarkable robustness in generating realistic images, even in cases where key input conditions—such as text or resolution—are missing. For instance, the model can generate forest images at various scales when provided with text like “There is a dense forest” without a specified resolution. Besides, when only the resolution is provided without specific textual descriptions, the model can generate resolution-specific images with diverse scenes such as a forest or an urban area. This ability highlights the model’s robustness in handling incomplete or missing input data, making it adaptable for real-world applications where input conditions may be partial. This ability benefits from our proposed dynamic condition adaptation strategy described in Section IV-B.
In Fig. 10, we tested whether the model could generate images with different levels of spatial detail given the same textual description but different resolution conditions. For example, using the prompt “Some white storage tanks are in a piece of bare land,” we generated images at 0.5m, 1m, and 2m per pixel resolutions. The resulting images exhibited variations in the relative size and density of the storage tanks that corresponded well with the specified resolutions, effectively mimicking real-world scale variations. Similarly, for the prompt “Many green trees are in a piece of forest,” the level of detail in tree structures varied appropriately with the resolution, demonstrating the model’s capability to produce resolution-dependent images.
@@ V-E Remote Sensing Image Editing
Importantly, Text2Earth ensures that these modifications are seamlessly integrated with the surrounding areas, maintaining continuity and coherence. This makes it an ideal tool for customized remote sensing image editing, catering to diverse applications such as urban planning.
@@ V-G1 Text-Driven Multi-Modal Image Generation
We fine-tuned Text2Earth using LoRA on the RSICD dataset to facilitate text-driven multi-modal image generation tasks, such as Text2SAR, Text2NIR, and Text2PAN. The results are illustrated in Fig. 14. The experiments demonstrate that Text2Earth can generate multi-modal images with high quality and semantic consistency. For instance, in the generated NIR images, green vegetation areas exhibit high pixel values, aligning with the physical imaging principles of NIR, where green vegetation reflects strongly in the near-infrared spectrum. These results underscore Text2Earth’s ability to effectively transfer its general knowledge to multi-modal remote sensing image generation tasks.
@@
Table IV shows quantitative evaluation on text-driven multi-modal image generation. FID scores across different modalities are not directly comparable because the FID for each modality is computed using an Inception V3 model pre-trained on data specific to that modality. Besides, unlike optical images (RGB, NIR, and PAN), SAR images are captured through microwave radar signals, which makes their visual appearance fundamentally different from optical images. SAR images often contain speckle noise and lack significant color and detailed texture information. These factors reduce the amount of semantic information available for scene classification tasks, rendering scene classification on SAR images inherently more challenging than on RGB, NIR, or PAN images. This leads to a low Zero-Shot Cls-OA score for the Text2SAR generation. In summary, the primary purpose of Table IV is to provide baseline benchmark results for future text-driven multi-modal image generation research rather than to directly compare performance across modalities.
@@ V-G2 Image-to-Image Translation
V-H More Applications
V-H1 Data Augmentation
We explored using Text2Earth-generated synthetic images as a data augmentation tool for remote sensing scene classification. Specifically, we selected 1,027 text-image-category triplets from the RSICD dataset as training samples and generated over 20,000 synthetic images based on their textual descriptions. We then trained four widely used classification models, including VGG-19, ResNet-18, ViT-B-16, and Swin-S, under two configurations: (i) using only the original training samples and (ii) using a combination of the original samples with the synthetic images. As demonstrated in Table V, all models showed consistent performance improvements when augmented with the generated images, thereby confirming that Text2Earth can serve as an effective data augmentation engine.
TABLE V: Accuracy of the downstream remote sensing image classification task w/ and w/o data augmentation.
Training Data VGG-19 ResNet-18 ViT-B-16 Swin-s
w./o. Augment 85.39 87.24 90.94 92.21
w. Augment 91.04 93.86 94.74 96.10
V-H2 Remote Sensing Vision-Language Contrastive Pretraining Foundation Model
We further explored the application of our Git-10M dataset to pretrain a vision-language foundation model using the contrastive learning framework. We named this model Git-RSCLIP. We then conducted zero-shot classification experiments on multiple publicly available remote sensing image classification datasets by computing the similarities of images and textualized scene category prompts to evaluate the performance of our Git-RSCLIP model. In Table VI, the experimental results demonstrate that Git-RSCLIP significantly outperforms previous remote sensing CLIP models, such as RemoteCLIP and GeoRSCLIP, confirming the effectiveness of our large-scale dataset. We have made the Git-RSCLIP model publicly available on our project page: https://github.com/Chen-Yang-Liu/Text2Earth
TABLE VI: Comparison of zero-Shot classification accuracy between our model and the previous CLIP Models on multiple remote sensing scene classification datasets.
Method OPTIMAL31 RSC11 RSICB128 WHURS19 RSSCN7 CLRS Average
CLIP [20] 60.00 45.29 25.23 77.41 52.25 56.48 52.78
SkyCLIP50 [79] 77.31 60.47 38.60 78.31 55.07 61.03 61.80
RemoteCLIP [87] 81.99 67.05 34.25 92.54 51.71 66.04 65.60
GeoRSCLIP [23] 83.33 67.37 35.48 89.45 62.54 69.67 67.97
Git-RSCLIP (Ours) 95.00 66.96 52.25 93.93 63.50 65.18 72.80
VI Limitation and Discussion for Text2Earth Model
While our Text2Earth model demonstrates robust performance in large-scale text-driven remote sensing image generation, it exhibits certain limitations that merit further discussion. One notable limitation is its inability to precisely control the number of objects specified in the textual descriptions, particularly when a large quantity is involved. For instance, as illustrated in Fig. 16, when given the prompt “Twelve storage tanks are near some green trees and buildings,” the model generated only nine storage tanks. Similarly, the model is asked to generate seven farmlands but generates eight in the last example. These results suggest that, although Text2Earth can capture numerical cues to a certain extent, it struggles with fine-grained numerical control—a capability that is critical for accurately reflecting detailed quantitative information in generated scenes.
This limitation likely arises from the inherent challenges of aligning textual numerical information with spatial visual content during the generative process. The current model primarily focuses on learning high-level semantic relationships rather than enforcing strict quantitative constraints on object counts. To address this issue, future research could explore the integration of specialized numerical reasoning modules or enhanced conditioning strategies that explicitly account for numerical details. Additionally, incorporating more training examples with explicit numerical descriptions may further improve the model’s ability to precisely control object quantity.
In summary, while Text2Earth represents a significant advancement in remote sensing image generation, addressing its limitations in object quantity control is an important avenue for future work. Improving this aspect will not only enhance the fidelity of generated images but also expand the model’s applicability in real-world remote sensing applications—such as urban planning and disaster assessment—where precise quantitative control is essential.
Figure 16: Some failure cases about inaccurate control over object quantity. These results suggest that, although Text2Earth can capture numerical cues to a certain extent, it struggles with fine-grained numerical control.
VII Future Work